OpenStudy (anonymous):

A football is kicked at an 50 degree angle to the horizontal , travels a horizontal distance of 20m before hitting the ground. what is the initial speed? please tell me how you would do this

OpenStudy (anonymous):

Range= u cos x *t , t= 2u sinx /g . => range = u^2 sin 2x/g 20=u^2 sin(100)/10 , u^2= 200/(sin(100)) = 203 u= 14.24 m/sec.